a bungee jumper leaps off a cliff 122 meters from the ground. her cord, when fully stretched, is 72 meters long, and it takes 6 seconds after the jump for it to extend fully. the distance between the jumper and the ground as a function of time can be modeled as a quadratic function. what is the equation represents her distance from the ground as a function of time?
Answer by stanbon(75887)

a bungee jumper leaps off a cliff 122 meters from the ground. her cord, when fully stretched, is 72 meters long, and it takes 6 seconds after the jump for it to extend fully. the distance between the jumper and the ground as a function of time can be modeled as a quadratic function. what is the equation represents her distance from the ground as a function of time?
----
h(t) = -4.9t^2+vot+ho
----
Note: 122-72 = 50 ft. (lowest point for the bungee)
50 = -4.9(6)^2 + vo*6 + 122
-----
-72 = -4.9*36 + 6*vo
-----
6*vo = 104.4
vo = 17.4
------
h(t) = -4.9t^2 + 17.4t + 122
--------------------------------
Cheers,
Stan H.
-------------